Thirteen more than five times a number is equal to the difference of 101 and three times the number.Find the number.
Answer by josgarithmetic(39617)

13+5n=101-3n, temporarily ignoring absolute value of the difference;
13+5n+3n=101
8n=101-13
8n=88
n=11.